OpenStudy (anonymous):

Last one I am stuck on. The number (1-I) is a root of the equation x^3-8x^2+14x-12=0. Find the other roots of this equation.

ganeshie8 (ganeshie8):

hint : complex roots always come in conjugate pairs

ganeshie8 (ganeshie8):

So, if \((1-i)\) is one root, its conjugate pair : \((1+i)\), will also be a rot

ganeshie8 (ganeshie8):

*root

OpenStudy (anonymous):

Should I group and factor (x^3 - 8x^2) and (14x -12)?

ganeshie8 (ganeshie8):

are u given options for this q ?

OpenStudy (anonymous):

No, any convent method.

OpenStudy (anonymous):

Convenient

ganeshie8 (ganeshie8):

you already have two roots : \(1+i\) \(1-i\)

ganeshie8 (ganeshie8):

maybe u can use that info to find the third root ?

ganeshie8 (ganeshie8):

familiar with long division ?

ganeshie8 (ganeshie8):

or, if u know rational root theorem, u can test and see if any factors of 12 work as roots

OpenStudy (anonymous):

Do I divide by -1?

ganeshie8 (ganeshie8):

u familiar wid rational root theorem ?

OpenStudy (anonymous):

We briefly went over in class, not very sure of my self

ganeshie8 (ganeshie8):

okay its simple actually rational root theorem says this : ``` rationals roots will be of form p/q where, p = factors of constant term q = factors of leading coefficient ```

ganeshie8 (ganeshie8):

your constant term = 12 leading coefficient = 1

ganeshie8 (ganeshie8):

So, u can test and see if u get luck wid any factors of 12 : 1, 2, 3, 4, 6, 12

ganeshie8 (ganeshie8):

you will find that "6" is the third root

ganeshie8 (ganeshie8):

since this is a cubic, there will be only 3 roots.

OpenStudy (anonymous):

I'm going to work the problem and I'll post what I get

ganeshie8 (ganeshie8):

okay

OpenStudy (anonymous):

Does this look right? (X-6)(x^2-2x+2)=0 X-6=0 X=6 (X^2-2x+2) (X^2-2x)=-2 (X^2-2x+1)=-1 (X-1)^2=-1 X='1-i, x=1+i,

ganeshie8 (ganeshie8):

yes, but u need to show work on how you got x = 6 as a root

OpenStudy (anonymous):

That would just be adding. 6 to get x by its self